OpenStudy (anonymous):

recipricoal of -5over9?

OpenStudy (anonymous):

Reciprocal means flip (-5/9) -> (-9/5) (-9/5)

OpenStudy (anonymous):

-9/5

OpenStudy (anonymous):

reciprocal means any number to exponent -1 hence flipping it

OpenStudy (saifoo.khan):

\[1\div \frac{-5}{9}\]

OpenStudy (saifoo.khan):

\[1 \times \frac{9}{-5}\]

OpenStudy (saifoo.khan):

therefore, 9/-5